The Pawlson Company's year-end balance sheet is shown below. Its cost of common equity is 16%, its before-tax cost of debt is 12%, and its marginal tax rate is 40%. Assume that the firm's long-term debt sells at par value. The firm’s total debt, which is the sum of the company’s short-term debt and long-term debt, equals $1,161. The firm has 576 shares of common stock outstanding that sell for $4.00 per share.

Assets Liabilities And Equity
Cash $ 120 Accounts payable and accruals $ 10
Accounts receivable 240 Short-term debt 51
Inventories 360 Long-term debt 1,110
Plant and equipment, net 2,160 Common equity 1,709
Total assets $2,880 Total liabilities and equity $2,880

Calculate Pawlson's WACC using market-value weights. Round your answer to two decimal places. Do not round your intermediate calculations.
%
Answer #1

Weight of Debt = 1,161/(1,161 + 2,304) = 0.3351

Weight of Equity = 1 - 0.3351 = 0.6649

WACC = 0.3351(1 - 0.40)(0.12) + 0.6649(0.16)

WACC = 13.05%